A sphere of mass 12kg and another one of mass 8 kg are moving towards each other in a smooth linear groove. The speed of the heavier sphere is 3 times that of the lighter sphere . After collision the heavier sphere has a speed of 4/1/6 m/s and the lighter sphere a speed of 6m/s . Calculate the speeds of the sphere before collisions


1
Expert's answer
2021-05-19T17:03:54-0400

Solution.

"m_1=12kg;"

"m_2=8kg;"

"v_{01}=3v;"

"v_{02}=v;"

"v_1=4\dfrac{1}{6}m/s;"

"v_2=6m\/s;"

"m_1\overrightarrow{v_{01}}+m_2\overrightarrow{v_{02}}=m_1\overrightarrow{v_{1}}+m_2\overrightarrow{v_{2}};"

"m_1v_{01}-m_2v_{02}=m_1v_1+m_2v_2;"

"12\sdot3v-8v=12\sdot4\dfrac{1}{6}+8\sdot6;"

"36v-8v=50+48;"

"26v=98;"

"v=98\/26;"

"v=3\dfrac{10}{13};"

"v_{01}=3\sdot3\dfrac{10}{13}=11\dfrac{4}{13}m/s;"

"v_{02}=3\dfrac{10}{13}m/s;"

Answer: "v_{01}=11\dfrac{4}{13}m/s;v_{02}=3\dfrac{10}{13}m/s."